# KernelDNA: Dynamic Kernel Sharing via Decoupled Naive Adapters

### v1

@@ 5 Experiments
In this section, to validate the effectiveness of our proposed KernelDNA, we primarily evaluate image classification accuracy on the ImageNet-1K [22] dataset and downstream task performance (object detection and image segmentation) on the MS-COCO [18] dataset. For simplicity, in the following tables, “KW” denotes the KernelWarehouse method, while “KDNA” represents our proposed KernelDNA. All experiments were conducted on a system equipped with an Nvidia GeForce RTX 4090 GPU and an AMD EPYC 7H12 64-Core Processor CPU. To ensure consistency, latency is reported with an input batch size of 1 and single-threaded CPU execution, while throughput is measured with a batch size of 128. To minimize variance, all results are averaged over 1,000 iterations.
@@ 5.1 Image Classification on ImageNet-1K
Experiment Setup: To benchmark against SOTA methods, we evaluate our approach on three representative architectures: standard ResNet18/50 [8], advanced ConvNeXt-Tiny [21], and compact MobileNetV2 [23]. All experiments are implemented using publicly available PyTorch official codebases222We can download related codes from the link., with training strategies aligned with those of KernelWarehouse for fair comparison. In our experiments, b denotes the parameter budget ratio of dynamic convolution methods relative to standard convolution. Notably, ConvNeXt-Tiny employs enhanced training strategies (e.g., advanced data augmentation and learning rate schedules), while other architectures follow conventional training strategies. Detailed configurations include: Optimization is SGD with momentum (0.9) and weight decay (1e-4/4e-5) for ResNet/MobileNetV2, and AdamW with weight decay (5e-2) for ConvNeXt-Tiny. The learning rate is initialized at 0.1 for ResNet, 5e-2 for MobileNetV2, and 4e-3 for ConvNeXt-Tiny, decayed via cosine annealing. Batch size is 256 for ResNet and MobileNetV2, and 1024 for ConvNeXt-Tiny. Training epochs are set to 100/150 for ResNet/MobileNetV2, and 300 for ConvNeXt-Tiny. This setup ensures reproducibility and fairness in comparing parameter efficiency and computational overhead across methods.
@@
Experiment Result: As can be seen from Tab. 1. KernelDNA (KDNA) achieves state-of-the-art accuracy-efficiency trade-offs across diverse architectures (ResNet18/50, MobileNetV2, ConvNeXt-Tiny) while maintaining hardware-friendly inference. Compared to dynamic convolution variants like DY-Conv and ODConv, KDNA reduces parameters by 1.2–5 (e.g., 9.24M vs. 44.90M on ResNet18) with comparable or lower FLOPs (2.28G vs. 1.82G for standard ResNet18), yet outperforms them in accuracy (e.g., 74.23% vs. 73.97% Top-1 on ResNet18). Unlike parameter-heavy methods such as CondConv (81M params), KDNA preserves near-original throughput (7,683 fps vs. ResNet18’s 8,463 fps) and minimizes CPU latency (28.43ms vs. ODConv’s 54.34ms).
| Models | Params | FLOPs | Throughput GPU(fps) | Latency CPU(ms) | Top-1 Acc(%) | Top-5 Acc(%) |
| ResNet18 | 11.69M | 1.82G | 8463 | 22.16 | 70.25 | 89.38 |
| + SE | 11.78M | 1.82G | 8519 | 23.68 | 70.98 | 90.03 |
| + CBAM | 11.78M | 1.82G | 6023 | 21.25 | 71.01 | 89.85 |
| + ECA | 11.69M | 1.82G | 8091 | 23.54 | 70.60 | 89.68 |
| + CGC | 11.69M | 2.19G | 1537 | 78.92 | 71.60 | 90.35 |
| + DCD | 14.70M | 1.90G | 5305 | 41.43 | 72.33 | 90.65 |
| + CondConv | 81.35M | 1.82G | 280 | 89.90 | 71.99 | 90.27 |
| + DY-Conv | 45.47M | 1.87G | 3308 | 46.15 | 72.76 | 90.79 |
| + ODConv | 44.90M | 1.82G | 1267 | 54.34 | 73.97 | 91.35 |
| + KW | 11.93M | 2.23G | 2272 | 69.62 | 73.67 | 91.17 |
| + KDNA | 9.24M | 2.28G | 7683 | 28.43 | 74.23 | 92.38 |
| ResNet50 | 25.56M | 4.11G | 2432 | 90.77 | 78.44 | 94.24 |
| + DY-Conv | 100.88M | 4.22G | 1161 | 155.63 | 79.00 | 94.27 |
| + ODConv | 90.67M | 4.12G | 560 | 132.39 | 80.62 | 95.16 |
| + KW | 28.05M | 6.15G | 922 | 138.78 | 80.38 | 95.19 |
| + KDNA | 22.78M | 4.33G | 2346 | 93.12 | 80.95 | 95.54 |
| MobiV2() | 1.97M | 0.10G | 12787 | 10.04 | 64.30 | 85.21 |
| + DY-Conv | 4.57M | 0.11G | 5818 | 45.59 | 69.05 | 88.37 |
| + ODConv | 4.44M | 0.11G | 3761 | 62.07 | 70.01 | 89.01 |
| + KW | 2.85M | 0.16G | 3284 | 26.43 | 68.29 | 87.93 |
| + KDNA | 1.88M | 0.11G | 12461 | 25.02 | 69.32 | 89.04 |
| MobiV2() | 3.50M | 0.31G | 6323 | 35.94 | 72.02 | 90.43 |
| + DY-Conv | 12.40M | 0.33G | 3667 | 46.93 | 74.94 | 91.83 |
| + ODConv | 11.52M | 0.32G | 2226 | 59.51 | 75.42 | 92.18 |
| + KW | 5.17M | 0.47G | 2288 | 44.58 | 74.68 | 91.90 |
| + KDNA | 3.14M | 0.31G | 6260 | 35.33 | 75.51 | 92.27 |
| ConvNeXt-Tiny | 28.59M | 4.47G | 1847 | 56.20 | 82.07 | 95.86 |
| + KW | 32.99M | 5.61G | 752 | 224.98 | 82.55 | 96.08 |
| + KDNA | 24.32M | 4.51G | 1739 | 60.12 | 82.76 | 96.34 |
@@
KDNA also surpasses semi-filters sharing approaches like KernelWarehouse (KW) in both efficiency and performance. With 1 parameter budgets, KDNA achieves higher accuracy (75.51% vs. KW’s 74.68% on MobileNetV2) while retaining 98% of the base model’s throughput (6,260 fps vs. 6,323 fps). Its static-dynamic decoupling ensures minimal computational overhead, outperforming KW’s kernel assembly (4.33G FLOPs vs. KW’s 6.15G on ResNet50). These results validate KDNA’s ability to unify parameter efficiency, hardware compatibility, and adaptive feature learning, making it a practical solution for real-world deployment.
@@ 5.2 Detection and Segmentation on MS-COCO
Experiment Setup: To further validate the generalization capability of our method, we conduct comparative experiments on downstream tasks, including object detection and instance segmentation, against state-of-the-art dynamic convolution approaches. Specifically, we evaluate on the MS-COCO dataset (118K training and 5K validation images), using Mask R-CNN [9] with ImageNet-1K pre-trained models as the backbone. For fairness, all models adhere to the same training strategies as KernelWarehouse [14], including identical data preprocessing (e.g., random resizing, flipping) and hyperparameter configurations. The evaluation metrics include Average Precision (AP) and its variants (AP50, AP75, APS, APM, APL) for object detection and instance segmentation. All experiments are conducted using the MMDetection [3] framework with the same evaluation settings.
@@
Experiment Result: Tab. 2 compares the performance of our proposed KernelDNA (KDNA) with state-of-the-art dynamic convolution methods on object detection and instance segmentation tasks. Evaluated across diverse backbones (ResNet50, MobileNetV2, ConvNeXt-Tiny), KDNA consistently outperforms existing approaches in multiple metrics, demonstrating superior parameter efficiency.
| Detectors | Backbone Models | Object Detection | Instance Segmentation | ||||||||||
| Mask R-CNN | ResNet50 | 39.6 | 61.6 | 43.3 | 24.4 | 43.7 | 50.0 | 36.4 | 58.7 | 38.6 | 20.4 | 40.4 | 48.4 |
| + DY-Conv | 39.6 | 62.1 | 43.1 | 24.7 | 43.3 | 50.5 | 36.6 | 59.1 | 38.6 | 20.9 | 40.2 | 49.1 | |
| + ODConv | 42.1 | 65.1 | 46.1 | 27.2 | 46.1 | 53.9 | 38.6 | 61.6 | 41.4 | 23.1 | 42.3 | 52.0 | |
| + KW | 41.8 | 64.5 | 45.9 | 26.6 | 45.5 | 53.0 | 38.4 | 61.4 | 41.2 | 22.2 | 42.0 | 51.6 | |
| + KDAN | 42.5 | 66.2 | 46.5 | 28.0 | 46.5 | 54.5 | 39.1 | 61.9 | 41.5 | 23.5 | 42.8 | 52.6 | |
| MobileNetV2 () | 33.8 | 55.2 | 35.8 | 19.7 | 36.5 | 44.4 | 31.7 | 52.4 | 33.3 | 16.4 | 34.4 | 43.7 | |
| + DY-Conv | 37.0 | 58.6 | 40.3 | 21.9 | 40.1 | 47.9 | 34.1 | 55.7 | 36.1 | 18.6 | 37.1 | 46.3 | |
| + ODConv | 37.2 | 59.4 | 39.9 | 22.6 | 40.0 | 48.0 | 34.5 | 56.4 | 36.3 | 19.3 | 37.3 | 46.8 | |
| + KW | 36.4 | 58.3 | 39.2 | 22.0 | 39.6 | 47.0 | 33.7 | 55.1 | 35.7 | 18.9 | 36.7 | 45.6 | |
| + KDAN | 37.3 | 59.8 | 40.5 | 23.0 | 40.7 | 48.5 | 34.8 | 56.5 | 36.8 | 19.5 | 37.8 | 47.2 | |
| ConvNeXt-Tiny | 43.4 | 65.8 | 47.7 | 27.6 | 46.8 | 55.9 | 39.7 | 62.6 | 42.4 | 23.1 | 43.1 | 53.7 | |
| + KW | 44.8 | 67.7 | 48.9 | 29.8 | 48.3 | 57.3 | 40.6 | 64.4 | 43.4 | 24.7 | 44.1 | 54.8 | |
| + KDAN | 45.0 | 67.9 | 49.2 | 30.2 | 48.7 | 58.0 | 41.0 | 65.1 | 43.6 | 24.8 | 44.5 | 55.2 | |
@@
Notably, KDNA surpasses 4-parameter methods like ODConv (detection : 42.5 vs. 42.1 on ResNet50) and outperforms KernelWarehouse (KW) under equal parameter constraints (detection : 37.3 vs. 36.4 on MobileNetV2). Its decoupled dynamic-static design enhances multi-scale adaptability, improving small-object detection () by +0.8 and large-object segmentation () by +0.4 over KW on ConvNeXt-Tiny. Crucially, KDNA maintains hardware efficiency via static attention pre-fusion, exceeding the throughput of 4-parameter methods while enabling plug-and-play deployment. This demonstrates its superior balance of accuracy, parameter efficiency, and practicality for real-world vision tasks.
@@ 5.3 Ablation Study
Settings of “parent” and “child” convolutional layers. Although previous experiments have demonstrated the effectiveness of our proposed dynamic kernel-sharing mechanism, the optimal number and placement of shared kernels (“child”) and base kernels (“parent”) require further exploration. Let the shared kernels be denoted as “S” and the base kernels (full kernels) as “F”. Subscripts of “S” and “F” indicate parent-child relationships. Considering that convolutional layers within the same block (typically with consistent kernel sizes) are more amenable to parameter sharing, our experiments focus on sharing and extending kernels across all blocks within the same stage. The results are shown in Tab. 3.
| Models | Parent-Child | Params | Throughput GPU(fps) | Top-1 Acc(%) | Top-5 Acc(%) |
| ResNet18 | FF-FF(original) | 11.69M | 8463 | 70.25 | 89.38 |
| FF-FFF | 14.64M | 7944 | 73.41 | 91.37 | |
| FF-FF-FF | 17.59M | 7496 | 74.07 | 92.93 | |
| FS-FSF | 9.07M | 7814 | 73.95 | 92.61 | |
| FS-SSF | 6.29M | 7756 | 73.42 | 91.64 | |
@@
From the results, it is evident that the placement of parent and child kernels is crucial for model performance. Using preceding layers as child kernels for a parent layer yields better results than using subsequent layers. Additionally, as the number of shared child kernels increases, the model’s performance gains diminish and may even degrade. Thus, simply scaling up the number of child kernels has an upper limit. However, scaling up the number of parent kernels gradually improves performance, albeit at the cost of increased parameters. Based on these experiments, we find that adopting the “FS-SF-SF” sequence and quantity within the same block achieves the best trade-off between inference efficiency and performance. The “-” indicates the position for the residual summation.
@@
Expanding Parent Kernels: Adding more static convolutional layers.
Direct Copying: Creating shared child kernels by directly duplicating parent kernels.
Adapter-Based Generation: Producing shared child kernels through lightweight adapters (our proposed method).
| Models | Increase layers | Params | Throughput GPU(fps) | Top-1 Acc(%) | Top-5 Acc(%) |
| ResNet18 | Original | 11.69M | 8463 | 70.25 | 89.38 |
| Expanding Parent | 14.64M | 7944 | 73.41 | 91.37 | |
| Direct Copying | 8.74M | 7460 | 69.85 | 89.41 | |
| Adapter-Based | 9.24M | 7683 | 74.23 | 92.38 | |
| ResNet50 | Original | 25.56M | 2432 | 78.44 | 94.24 |
| Expanding Parent | 26.67M | 2361 | 80.23 | 94.83 | |
| Direct Copying | 22.55M | 2355 | 75.21 | 92.87 | |
| Adapter-Based | 22.78M | 2346 | 80.95 | 95.54 |
The results are presented in Tab. 4. From the experimental outcomes, it is evident that our adapter-based approach achieves a superior trade-off between performance and efficiency compared to the other two methods.
Three attention components analysis. Our comparative study on different methods of expanding convolutional layers demonstrates that sharing parameters via a “bridge” (i.e., the adapter) is critical, yet the design of this bridge is equally pivotal. In this work, we decouple visual attention mechanisms across dimensions into dynamic and static components, addressing channel-wise, spatial-wise, and filter-wise kernel modulation. Compared to prior works like ODConv, which employs additional dynamic and sequential kernel attention, our implementation is simpler and more efficient.
| Models | Channel | Spatial | Filter | Params | Throughput GPU(fps) | Top-1 Acc(%) | Top-5 Acc(%) |
| ResNet18 | ✗ | ✗ | ✗ | 8.74M | 7460 | 69.85 | 89.41 |
| ✓ | ✗ | ✗ | 9.24M | 7695 | 71.82 | 90.72 | |
| ✗ | ✓ | ✗ | 8.74M | 7732 | 70.90 | 90.06 | |
| ✗ | ✗ | ✓ | 8.75M | 7786 | 70.29 | 89.29 | |
| ✓ | ✓ | ✗ | 9.24M | 7485 | 72.68 | 91.54 | |
| ✓ | ✗ | ✓ | 9.24M | 7499 | 71.98 | 91.05 | |
| ✗ | ✓ | ✓ | 8.75M | 7606 | 71.01 | 90.23 | |
@@
To further validate the effectiveness of the three attention components, we conduct ablation studies (see Tab. 5). The results show that all three attention mechanisms contribute positively to performance. Notably, Channel Attention delivers the most significant accuracy gains, albeit with a slight impact on inference efficiency. In contrast, Spatial Attention and Filter Attention achieve moderate performance improvements without introducing additional computational overhead. These findings collectively confirm the efficacy of our lightweight adapter’s tripartite attention design.

### v2

@@ 5 Experiments
In this section, to validate the effectiveness of our proposed KernelDNA, we primarily evaluate image classification accuracy on the ImageNet-1K [russakovsky2015imagenet] dataset and downstream task performance (object detection and image segmentation) on the MS-COCO [lin2014microsoft] dataset. For simplicity, in the following tables, “KW” denotes the KernelWarehouse method, while “KDNA” represents our proposed KernelDNA. All experiments were conducted on a system equipped with an Nvidia GeForce RTX 4090 GPU and an AMD EPYC 7H12 64-Core Processor CPU. To ensure consistency, latency is reported with an input batch size of 1 and single-threaded CPU execution, while throughput is measured with a batch size of 128. To minimize variance, all results are averaged over 1,000 iterations. The models are trained from scratch, consistent with all the other compared methods.
@@ 5.1 Image Classification on ImageNet-1K
Experiment Setup: To benchmark against SOTA methods, we evaluate our approach on three representative architectures: standard ResNet18/50 [he2016deep], advanced ConvNeXt-Tiny [liu2022convnet], and compact MobileNetV2 [sandler2018mobilenetv2]. All experiments are implemented using publicly available PyTorch official codebases, with training strategies aligned with those of KernelWarehouse for fair comparison. Notably, ConvNeXt-Tiny employs enhanced training strategies (e.g., advanced data augmentation and learning rate schedules), while other architectures follow conventional training strategies. Detailed configurations include: optimizer is SGD with momentum (0.9) and weight decay (1e-4/4e-5) for ResNet/MobileNetV2, and AdamW with weight decay (5e-2) for ConvNeXt-Tiny. The learning rate is initialized at 0.1 for ResNet, 5e-2 for MobileNetV2, and 4e-3 for ConvNeXt-Tiny, decayed via cosine annealing. Batch size is 256 for ResNet and MobileNetV2, and 1024 for ConvNeXt-Tiny. Training epochs are set to 100/150 for ResNet/MobileNetV2, and 300 for ConvNeXt-Tiny. This setup ensures reproducibility and fairness in comparing parameter efficiency and computational overhead across methods.
@@
Experiment Result: As can be seen from Table 1, KDNA achieves SOTA accuracy-efficiency trade-offs across diverse architectures (ResNet18/50, MobileNetV2, ConvNeXt-Tiny) while maintaining hardware-friendly inference. Compared to dynamic convolution variants like DY-Conv and ODConv, KDNA reduces parameters by 1.2–5 (e.g., 9.24M vs. 44.90M on ResNet18) with comparable or lower FLOPs, yet outperforms them in accuracy (e.g., 74.23% vs. 73.97% Top-1 on ResNet18). Unlike parameter-heavy methods such as CondConv (81M params), KDNA preserves near-original throughput (7,683 fps vs. ResNet18’s 8,463 fps) and minimizes CPU latency (28.43ms vs. ODConv’s 54.34ms). FDConv achieves comparable parameter efficiency (11.75M vs. 9.24M on ResNet18) but suffers from severe inference overhead (422 fps, 18 slower than KDNA) with inferior accuracy (71.25% vs. 74.23% Top-1), demonstrating that frequency-domain transformations introduce prohibitive computational costs.
| Model | Pm (M) | FP (G) | TP GPU (fps) | Lt CPU (ms) | Top-1 Acc(%) | Top-5 Acc(%) |
| ResNet18 | 11.69 | 1.82 | 8463 | 22.16 | 70.25 | 89.38 |
| + SE | 11.78 | 1.82 | 8319 | 23.68 | 70.98 | 90.03 |
| + CBAM | 11.78 | 1.82 | 6023 | 25.25 | 71.01 | 89.85 |
| + ECA | 11.69 | 1.82 | 8372 | 23.54 | 70.60 | 89.68 |
| + CGC | 11.69 | 2.19 | 1537 | 78.92 | 71.60 | 90.35 |
| + DCD | 14.70 | 1.90 | 5305 | 41.43 | 72.33 | 90.65 |
| + CondConv | 81.35 | 1.82 | 280 | 89.90 | 71.99 | 90.27 |
| + DY-Conv | 45.47 | 1.87 | 3308 | 46.15 | 72.76 | 90.79 |
| + ODConv | 44.90 | 1.82 | 1267 | 54.34 | 73.97 | 91.35 |
| + KW | 11.93 | 2.23 | 2272 | 69.62 | 73.67 | 91.17 |
| + FDConv | 11.75 | 1.92 | 422 | 1013 | 71.25 | 90.04 |
| + KDNA | 9.24 | 2.28 | 7683 | 28.43 | 74.23 | 92.38 |
| ResNet50 | 25.56 | 4.11 | 2432 | 90.77 | 78.44 | 94.24 |
| + DY-Conv | 100.88 | 4.22 | 1161 | 155.63 | 79.00 | 94.27 |
| + ODConv | 90.67 | 4.12 | 560 | 132.39 | 80.62 | 95.16 |
| + KW | 28.05 | 6.15 | 922 | 138.78 | 80.38 | 95.19 |
| + FDConv | 27.16 | 4.29 | 230 | 1380.65 | 78.76 | 94.35 |
| + KDNA | 22.78 | 4.33 | 2346 | 93.12 | 80.95 | 95.54 |
| MbV2() | 1.97 | 0.10 | 12787 | 10.04 | 64.30 | 85.21 |
| + DY-Conv | 4.57 | 0.11 | 5818 | 45.59 | 69.05 | 88.37 |
| + ODConv | 4.44 | 0.11 | 3761 | 62.07 | 70.01 | 89.01 |
| + KW | 2.85 | 0.16 | 3284 | 26.43 | 68.29 | 87.93 |
| + KDNA | 1.88 | 0.11 | 12461 | 10.27 | 69.32 | 89.04 |
| MbV2() | 3.50 | 0.31 | 6323 | 30.21 | 72.02 | 90.43 |
| + DY-Conv | 12.40 | 0.33 | 3667 | 46.93 | 74.94 | 91.83 |
| + ODConv | 11.52 | 0.32 | 2226 | 59.51 | 75.42 | 92.18 |
| + KW | 5.17 | 0.47 | 2288 | 44.58 | 74.68 | 91.90 |
| + KDNA | 3.14 | 0.31 | 6260 | 33.46 | 75.51 | 92.27 |
| Conv-T | 28.59 | 4.47 | 1847 | 56.20 | 82.07 | 95.86 |
| + KW | 32.99 | 5.61 | 752 | 224.98 | 82.55 | 96.08 |
| + KDNA | 24.32 | 4.51 | 1739 | 60.12 | 82.76 | 96.34 |
@@
KDNA also surpasses semi-filters sharing approaches like KernelWarehouse (KW) in both efficiency and performance. With 61% parameter budgets, KDNA achieves higher accuracy (75.51% vs. KW’s 74.68% on MobileNetV2()) while retaining 99% of the base model’s throughput (6,260 fps vs. 6,323 fps). Its static-dynamic decoupling ensures minimal computational overhead, outperforming KW’s kernel assembly (4.33G FLOPs vs. KW’s 6.15G FLOPs on ResNet50). On ResNet50, FDConv exhibits comparable parameter efficiency (27.16M vs. 22.78M) but suffers from severe inference slowdown (230 fps, 10.2 slower than KDNA) with inferior accuracy (78.76% vs. 80.95% Top-1), highlighting the importance of maintaining standard convolution’s efficient structure. These results validate KDNA’s ability to unify parameter efficiency, hardware compatibility, and adaptive feature learning, making it a practical solution for real-world deployment.
@@ 5.2 Detection and Segmentation on MS-COCO
Experiment Setup: To further validate the generalization capability of our method, we conduct comparative experiments on downstream tasks, including object detection and instance segmentation, against state-of-the-art dynamic convolution approaches. Specifically, we evaluate on the MS-COCO dataset (118K training and 5K validation images), using Mask R-CNN [he2017mask] with ImageNet-1K pre-trained models as the backbone. For fairness, all models adhere to the same training strategies as KernelWarehouse [Li2024], including identical data preprocessing (e.g., random resizing, flipping) and hyperparameter configurations. The evaluation metrics include Average Precision (AP) and its variants (AP50, AP75, APS, APM, APL) for object detection and instance segmentation. All experiments are conducted using the MMDetection [mmdetection] framework with the same evaluation settings.
@@
Experiment Result: Table 2 compares the performance of our proposed KernelDNA (KDNA) with state-of-the-art dynamic convolution methods on object detection and instance segmentation tasks. Evaluated across diverse backbones (ResNet50, MobileNetV2, ConvNeXt-Tiny), KDNA consistently outperforms existing approaches in multiple metrics, demonstrating superior parameter efficiency.
Object Detection | Instance Segmentation | ||||||||||||
ResNet50 | 39.6 | 61.6 | 43.3 | 24.4 | 43.7 | 50.0 | 36.4 | 58.7 | 38.6 | 20.4 | 40.4 | 48.4 | ||
| + DY-Conv | 39.6 | 62.1 | 43.1 | 24.7 | 43.3 | 50.5 | 36.6 | 59.1 | 38.6 | 20.9 | 40.2 | 49.1 | |||
| + ODConv | 42.1 | 65.1 | 46.1 | 27.2 | 46.1 | 53.9 | 38.6 | 61.6 | 41.4 | 23.1 | 42.3 | 52.0 | |||
| + KW | 41.8 | 64.5 | 45.9 | 26.6 | 45.5 | 53.0 | 38.4 | 61.4 | 41.2 | 22.2 | 42.0 | 51.6 | |||
| + KDNA | 42.5 | 66.2 | 46.5 | 28.0 | 46.5 | 54.5 | 39.1 | 61.9 | 41.5 | 23.5 | 42.8 | 52.6 | |||
| MobileNetV2 () | 33.8 | 55.2 | 35.8 | 19.7 | 36.5 | 44.4 | 31.7 | 52.4 | 33.3 | 16.4 | 34.4 | 43.7 | |||
| + DY-Conv | 37.0 | 58.6 | 40.3 | 21.9 | 40.1 | 47.9 | 34.1 | 55.7 | 36.1 | 18.6 | 37.1 | 46.3 | |||
| + ODConv | 37.2 | 59.4 | 39.9 | 22.6 | 40.0 | 48.0 | 34.5 | 56.4 | 36.3 | 19.3 | 37.3 | 46.8 | |||
| + KW | 36.4 | 58.3 | 39.2 | 22.0 | 39.6 | 47.0 | 33.7 | 55.1 | 35.7 | 18.9 | 36.7 | 45.6 | |||
| + KDNA | 37.3 | 59.8 | 40.5 | 23.0 | 40.7 | 48.5 | 34.8 | 56.5 | 36.8 | 19.5 | 37.8 | 47.2 | |||
| ConvNeXt-Tiny | 43.4 | 65.8 | 47.7 | 27.6 | 46.8 | 55.9 | 39.7 | 62.6 | 42.4 | 23.1 | 43.1 | 53.7 | |||
| + KW | 44.8 | 67.7 | 48.9 | 29.8 | 48.3 | 57.3 | 40.6 | 64.4 | 43.4 | 24.7 | 44.1 | 54.8 | |||
| + FDConv | 45.2 | 67.9 | 48.8 | 29.9 | 48.7 | 57.6 | 40.8 | 64.8 | 43.6 | 24.8 | 44.5 | 54.9 | |||
| + KDNA | 45.4 | 68.1 | 49.4 | 30.4 | 48.9 | 58.2 | 41.2 | 65.3 | 43.8 | 24.9 | 44.7 | 55.4 | |||
@@
Notably, KDNA surpasses 4-parameter methods like ODConv (detection : 42.5 vs. 42.1 on ResNet50) and outperforms KernelWarehouse (KW) under equal parameter constraints (detection : 37.3 vs. 36.4 on MobileNetV2). On ConvNeXt-Tiny, FDConv achieves competitive results (detection : 45.2, segmentation : 40.8), but KDNA consistently outperforms FDConv across all metrics (detection : 45.4 vs. 45.2, segmentation : 41.2 vs. 40.8), particularly excelling in small-object detection (: 30.4 vs. 29.9) and large-object segmentation (: 58.2 vs. 57.6). Crucially, KDNA maintains hardware efficiency via static attention pre-fusion, while FDConv’s frequency-domain operations incur significant computational overhead. The consistent superiority of KDNA across all backbones demonstrates its superior balance of accuracy, parameter efficiency, and practicality for real-world vision tasks.
@@ 5.3 Ablation Study
Settings of “parent” and “child” convolutional layers. Although previous experiments have demonstrated the effectiveness of our proposed dynamic kernel-sharing mechanism, the optimal number and placement of shared kernels (“child”) and base kernels (“parent”) require further exploration. Let the shared kernels be denoted as “S” and the base kernels (full kernels) as “F”. Subscripts of “S” and “F” indicate parent-child relationships. Considering that convolutional layers within the same stage (typically with consistent kernel sizes) are more amenable to parameter sharing, our experiments focus on sharing and extending kernels across all blocks within the same stage, as is shown in Table 3.
| Model | Parent-Child | Params |
FF-FF(orig) | 11.69M | 8463 | 70.25 | 89.38 | ||||||
| FF-FFF | 14.64M | 7944 | 73.41 | 91.37 | |||||||
| FF-FF-FF | 17.59M | 7496 | 74.07 | 92.23 | |||||||
| FS-FSF | 9.07M | 7814 | 73.95 | 92.61 | |||||||
| FS-SSF | 6.29M | 7756 | 73.42 | 91.64 | |||||||
@@
From the results, it is evident that the placement of parent and child kernels is crucial for model performance. Using preceding layers as child kernels for a parent layer yields better results than using subsequent layers. Additionally, as the number of shared child kernels increases, the model’s performance gains diminish and may even degrade. Thus, simply scaling up the number of child kernels has an upper limit. However, scaling up the number of parent kernels gradually improves performance, albeit at the cost of increased parameters. Based on these experiments, we find that adopting the “FS-SF-SF” sequence and quantity within the same block achieves the best trade-off between inference efficiency and performance. Note, the “-” indicates the position for the residual summation, “orig” indicates the original sequence.
@@
Expanding Parent Kernels (Expand): Adding more static convolutional layers.
Direct Copying (Copy): Creating shared child kernels by directly duplicating parent kernels.
Adapter-Based Generation (Adapter): Producing shared child kernels through lightweight adapters (Ours).
| Model |
Params |
Orig | 11.69M | 8463 | 70.25 | 89.38 | ||||||||
| Expand | 14.64M | 7744 | 73.41 | 91.37 | |||||||||
| Copy | 8.74M | 7735 | 69.85 | 89.41 | |||||||||
| Adapter | 9.24M | 7683 | 74.23 | 92.38 | |||||||||
Orig | 25.56M | 2432 | 78.44 | 94.24 | ||||||||
| Expand | 26.67M | 2358 | 80.23 | 94.83 | |||||||||
| Copy | 22.55M | 2355 | 75.21 | 92.87 | |||||||||
| Adapter | 22.78M | 2346 | 80.95 | 95.54 |
The results are presented in Table 4. From the experimental outcomes, it is evident that our adapter-based approach achieves a superior trade-off between performance and efficiency compared to the other two methods.
Analysis of the three attention components. Our comparative study on different methods of expanding convolutional layers highlights the importance of parameter sharing through a “bridge” (i.e., the adapter), and the design of this bridge is equally crucial. In this work, we decouple visual attention mechanisms across dimensions into dynamic and static components, addressing channel-wise, spatial-wise, and filter-wise kernel modulation. Compared to previous works like ODConv, which employs additional dynamic and sequential kernel attention, our implementation is simpler and more efficient.
| Channel | Spatial | Filter | Params | TP GPU (fps) | Top-1 Acc(%) | Top-5 Acc(%) |
| ✗ | ✗ | ✗ | 8.74M | 7863 | 69.85 | 89.41 |
| ✓ | ✗ | ✗ | 9.24M | 7726 | 71.82 | 90.72 |
| ✗ | ✓ | ✗ | 8.74M | 7815 | 70.90 | 90.06 |
| ✗ | ✗ | ✓ | 8.75M | 7839 | 70.29 | 89.29 |
| ✓ | ✓ | ✗ | 9.24M | 7698 | 72.68 | 91.54 |
| ✓ | ✗ | ✓ | 9.24M | 7699 | 71.98 | 91.05 |
| ✗ | ✓ | ✓ | 8.75M | 7804 | 71.01 | 90.23 |
@@
To further validate the effectiveness of the three attention components, we conduct ablation studies (see Table 5). The results show that all three attention mechanisms contribute positively to performance. Notably, Channel Attention delivers the most significant accuracy gains, albeit with a slight impact on inference efficiency. In contrast, Spatial Attention and Filter Attention achieve moderate performance improvements without introducing additional computational overhead. These findings collectively confirm the efficacy of our lightweight adapter’s tripartite attention design.
5.4 Parameter Scaling
KDNA also supports flexible parameter scaling by adjusting the number of parent and child kernels. Table 6 presents the results of scaling KDNA on both ResNet18 and MobileNetV2(). As the scale increases, KDNA consistently improves accuracy with a corresponding increase in parameter count, demonstrating its scalability across different model sizes.
| Model | Scale | Params |
| ResNet18 | baseline | 11.69M | 70.25 | 89.38 | ||||
| KDNA (1) | 9.24M | 74.23 | 92.38 | |||||
| KDNA (2) | 18.48M | 75.41 | 92.75 | |||||
| KDNA (4) | 45.91M | 76.58 | 93.26 | |||||
| MobileNetV2() | baseline | 1.97M | 64.30 | 85.21 | ||||
| KDNA (1) | 1.88M | 69.32 | 89.04 | |||||
| KDNA (2) | 3.24M | 70.82 | 89.73 | |||||
| KDNA (4) | 4.46M | 71.36 | 90.22 |